# Which Factors Make Code LLMs More Vulnerable to Backdoor Attacks? A Systematic Study

### v1

@@ IV-C RQ3: How Do Inference Factors Affect Backdoor Attacks?
Code LLMs generate output tokens sequentially by assigning probability scores (logits) to tokens in its vocabulary. Inference factors control how these logit scores guide token selection, leading to variations in the generated outputs. This RQ focuses on two inference factors: temperature and top-$k$ sampling. We select the poisoned models from Section IV-A1 under three poisoning rates (low:0.05%, medium:0.5%, high:5%). Results for 0.5% and 5% poisoning rates are not shown in Figure 10 as all ASRs stay above 94% (absolute difference $<$1%), showing that the logit score of the poisoned token is much higher than that of other tokens, limiting the influence of sampling strategies. Both temperature and top-$k$ sampling show no significant impact on FTR.
@@ IV-C1 Temperature Sampling
We hypothesize that higher $T$ can reduce ASR. Intuitively, increasing $T$ increases the likelihood of selecting non-poisoned tokens that have lower logit scores. Figure 10 shows the ASR results. Wilcoxon signed-rank tests confirm that under poisoning rates of 0.05% and 0.5%, ASR decreases as $T$ increases. Nearly all comparisons are Bonferroni significant (p$<$0.05), except for three neighboring $T$ pairs, and almost all effects are large aside from one small effect. At 0.05% poisoning rate, increasing $T$ from 0 to 1.2 consistently reduces ASR across all models and trigger types. For instance, shown in Figure 10, with CodeT5, this temperature increase causes fixed and grammar triggers to nearly halve their ASRs.
Finding 8: Higher temperature reduces ASR effectively at low poisoning rates. Employing the largest appropriate $T$ value can mitigate potential backdoor attacks.
@@ IV-C2 Top-$k$ Sampling
We experiment with $k$ values from 1-5 (step=1) and 10-50 (step=10). Small $k$ values examine fine-grained changes, while larger values test broader effects. Figure 10 shows the ASR results. Wilcoxon signed-rank test shows that the ASR under $k=1$ is significantly higher than under larger $k$ values (Bonferroni-corrected p$<$0.05, large effect). Other comparisons are not always Bonferroni significant. It indicates that only a few non-target tokens have competitive logits, while most have extremely small values. Increasing $k$ beyond 2 yields diminishing returns in reducing ASR.
@@
Finding 9: Higher top-$k$ values reduce ASR, with the most significant drop occurring from $k=1$ to $2$.
@@ V Discussion
V-A Case Study: Backdoor Defense Under Low Poisoning Rate
In the previous sections, we evaluate backdoor attacks under different factors. Our findings reveal a critical vulnerability overlooked in prior studies–backdoor attacks are effective even with extremely low poisoning rates and minimal poisoned samples–e.g., 8 in 10,000 in Section IV-A1 and 20 in 300,000 in Section IV-A3. The effectiveness of previous proposed defense methods under such low poisoning rates remains largely unexplored, motivating us to evaluate. For instance, Ramakrishnan et al. [4]’s evaluation of spectral signature focuses on poisoning rates above 1%.
Following Mu et al. [12], we identify 5 Code LLM backdoor defense methods widely used in prior studies by removing poisoned samples/trigger words from the dataset. Spectral signature [4] and activation clustering [39] remove poisoned samples by identifying distinct patterns in their feature representations and activation patterns that differentiate them from normal samples. CodePurify [12], ONION [13], and OSeql [25] identify trigger tokens by analyzing model behavior changes when masking potentially poisoned tokens or lines of code.
Ramakrishnan et al. [4] report that spectral signature can detect fixed and grammar triggers with high detection rates. Yang et al. [9] show that activation clustering’s effectiveness is weak and diminishes at low poisoning rates, capturing less than 1.24% of poisoned samples at 0.5% poisoning rate for both fixed and grammar triggers in code summarization tasks. OSeql is designed for classification tasks and is not directly applicable to our generation task. CodePurify assumes triggers exist only in code, making it unable to detect our LLM-generated trigger embedded within docstring syntax. ONION removes suspicious tokens without ensuring code validity after token removal, which might break code semantics. Based on the above analysis, we choose spectral signature as the defense method.
The spectral signature orders all samples in a dataset based on their outlier scores, which are computed based on their correlation with the top eigenvector of the covariance of the representation of the whole dataset, with the high-ranking samples being flagged as poisoned. Following Yang et al.’s implementation [9], CodeBERT [40] is used as encoder and a 6-layer Transformer decoder is employed. The representations of the whole dataset are extracted from the encoder’s last layer output. We rank all samples according to their outlier scores and identify the highest $\beta\times\texttt{Poisoning Rate}$ samples as poisoned samples, where $\beta$ represents the removal rate and $\beta=1.5$ is used following Yang et al. [9].
Two poisoned datasets are chosen. I: 8 poisoned samples in a total of 10,000 samples (0.08% poisoning rate). Detailed in Section IV-A1, CodeT5 achieves ASRs of 51.4%, 42.4%, and 71.1% for fixed, grammar, and LLM-generated triggers respectively. CodeT5+ shows $>$90% ASR across all triggers, while PLBART reaches 89.1% and 59.2% for fixed and grammar triggers. II: 20 poisoned samples in a total of 300,000 samples (0.0067% poisoning rate). Detailed in Section IV-A3, CodeT5 achieves an ASR of 82.9% and 51.8% for the fixed and grammar triggers, respectively. CodeT5+ achieves an ASR of 61.1%, 10.3% and 49.4% respectively. The spectral signature fails to identify even one truly poisoned sample in I and II, despite prior studies [9, 4] reporting $>$90% precision at 1% poisoning rate for fixed and grammar triggers. Furthermore, for II, it becomes computationally impractical due to its time complexity of $O(n^{2})$ introduced by matrix multiplication, where $n$ is the number of samples in the dataset. This highlights the urgent need for efficient and scalable defense methods that can remove a few poisoned samples within a large-scale dataset.
V-B Threats to Validity
@@
Threats to Internal Validity. We take several measures to minimize internal validity threats. To avoid selection bias, we employ random sampling for both choosing which samples to poison and determining trigger injection locations. Our fine-tuning leverages CodeT5+’s official script and loads pre-trained parameters from HuggingFace’s model hub [41]. The BLEU-4 scores for three models on clean data closely match those reported in their original papers. Thus, we believe that the threats to internal validity are minimal.
@@
Threats to External Validity. Our study examines code summarization, a representative code-text generation task, and evaluates three widely-used Code LLMs: CodeT5, CodeT5+, and PLBART, on a large-scale dataset, CodeSearchNet’s Java dataset. However, the Code LLM ecosystem is vast and diverse, potentially limiting our findings’ generalizability to other Code LLMs, tasks, and programming languages. Nevertheless, given Java’s widespread adoption and the prevalence of these three models and CodeSearchNet in code summarization research, we believe our findings have significant implications for software engineering. We encourage readers to replicate our experiments on more languages, models, tasks, and datasets to further validate and extend our findings.
@@
Threats to Construct Validity. While other metrics such as Exact Match [42] exist, we follow prior work on Code LLM backdoor attacks [9] and standard practice in code summarization [10] by using BLEU-4 to assess poisoned models’ performance degradation on clean data. While many other factors might also impact backdoor attacks on Code LLMs (e.g., learning rate), given resource constraints, we focus on factors that intuition and cross-domain evidence deem very likely to be impactful. Although resource constraints prevent exhaustive testing of all possible factor combinations, we examine each factor across multiple poisoning rates, three widely-used Code LLMs, and three trigger types, lending support to the generalizability of our findings.
@@ VI-A Backdoor Attacks: from Machine Learning to Code LLMs
Hussain et al. [26] conduct a survey of poisoning attacks on Code LLMs, proposing a taxonomy that systematically categorizes existing works, including pioneering work by Ramakrishnan et al. [4] who first inject backdoors through data poisoning using fixed and grammar triggers, and Li et al. [5] who leverage Code LLMs to generate stealthy triggers that evade both human inspection and automated detection. The above three trigger types are employed in our work. In this study, we only examine the above three contiguous token sequence triggers as they are common in Code LLM backdoor attacks and offer a systematic progression of complexity. Beyond a contiguous token sequence, there are other trigger forms. For example, Yang et al. [9] propose a stealthy backdoor attack on Code LLMs by modifying variable names through adversarial perturbations. Aghakhani et al. [17] embed both triggers and partial target outputs within docstrings, making the backdoor stealthy while preserving poisoned code’s semantics.
Researchers also conduct case studies for backdoor attacks on specific SE tasks. Sun et al. [6] stealthily boost buggy or vulnerable code into the top 11% of rankings by altering just one variable or function name. Schuster et al. [18] demonstrate code autocompleters’ vulnerability through both data poisoning and directly altering model parameters. Nguyen et al. [52] reveal that three state-of-the-art API recommender systems are vulnerable to backdoor attacks through data poisoning. Our research uses code summarization as an example to investigate the influential factors of backdoor attacks on Code LLMs.
@@ VI-B Defending Code LLMs against Data-Poisoning Backdoors
Removing suspicious samples is a common defense strategy. Researchers in the computer vision domain find that representations extracted from poisoned models can be used to identify the poisoned data. Tran et al. [53] use spectral signatures calculated using the output of model’s hidden layers to remove poisoned images. Ramakrishnan et al. [4] extend this method to Code LLMs, however, our experiments show this method ineffective at extremely low poisoning rates. Chen et al. [39] identify poisoned images by clustering neuron activations, leveraging the observation that poisoned inputs often exhibit distinct activation patterns compared to clean ones. Yang et al. [9] migrate this method to Code LLMs and find it ineffective.
There are also defenses aim to remove triggers from samples. Qi et al. [13] identify potential trigger words using perplexity. Hussain et al. [25] propose an occlusion-based method to identify trigger tokens by observing that removing trigger tokens significantly alters the model’s prediction confidence, while removing other code segments does not. Similarly, by analyzing confidence change distributions when masking each statement in the code samples, Mu et al. [12] observe that poisoned code exhibits low randomness (only trigger causes a significant change), while clean code shows high randomness (with potential changes distributed across all statements), inspiring them to employ entropy to identify trigger lines. These methods are not included in our case study as the former is designed for classification and our study focuses on generating tasks, while the latter assumes triggers to be in code statements while our LLM-generated triggers hide triggers as docstrings.
@@ VII Conclusion and Future Work
In this study, we investigate the impact of various factors on backdoor attacks targeting Code LLMs. Focusing on code summarization, we conduct extensive experiments using three trigger types—fixed, grammar, and LLM-generated—across three widely used Code LLMs under multiple data, training, and inference factors. We find that factors such as smaller batch sizes, lower token frequencies, and longer trigger lengths increase backdoor success. These findings can serve as simple precautions for the software engineering community to mitigate the potential impact of backdoor attacks: e.g., users of Code LLMs can increase sampling randomness (e.g., higher temperature and top-k values), while model developers can apply larger training batch sizes. Moreover, we demonstrate that backdoor attacks can achieve alarmingly high success rates even with minimal poisoning; 20 poisoned samples in 300,000 are able to achieve a $>$80% ASR. This shows the limitation of previous studies’ evaluations, which focus on higher poisoning rates, and calls for more effective countermeasures, as spectral signature is ineffective at cleaning such poisoned dataset with low poisoning rates. In future work, we plan to investigate other influential factors, and develop more effective and efficient defense methods that can remove sparse poisoned samples from large-scale training datasets.
Code, documentation, and full experiment results:
https://anonymous.4open.science/r/Backdoor-31C4

### v2

@@ IV-C RQ3: How Do Inference Factors Affect Backdoor Attacks?
Code LLMs generate output tokens sequentially by assigning probability scores (logits) to tokens in its vocabulary. Inference factors control how these logit scores guide token selection. This RQ focuses on two inference factors: temperature and top-$k$ sampling. We reuse the poisoned models from Section IV-A1 under three poisoning rates (low:0.05%, medium:0.5%, high:5%). Results for 0.5% and 5% poisoning rates are not shown in Figure 10 as all ASRs stay above 94% (absolute difference <<1%). Both temperature and top-$k$ sampling show no significant impact on FTR.
@@ IV-C1 Temperature Sampling
We hypothesize that higher $T$ can reduce ASR as it increases the likelihood of selecting non-poisoned tokens with lower logit scores. Figure 10 shows the ASR results. Wilcoxon signed-rank tests confirm that under poisoning rates of 0.05% and 0.5%, ASR decreases as $T$ increases. Nearly all comparisons are Bonferroni significant (p<<0.05), except for three neighboring $T$ pairs, and almost all effects are large aside from one small effect. At 0.05% poisoning rate, increasing $T$ from 0 to 1.2 consistently reduces ASR across all models and trigger types. Shown in Figure 10, with CodeT5, this temperature increase causes fixed and grammar triggers to nearly halve their ASRs.
Finding 8: Higher temperature reduces ASR effectively at low poisoning rates (0.05% and 0.5%). Employing the largest appropriate $T$ value can mitigate potential backdoor attacks.
@@ IV-C2 Top-$k$ Sampling
We experiment with $k$ values from 1-5 (step=1) and 10-50 (step=10). Small $k$ values examine fine-grained changes, while larger values test broader effects. Figure 10 shows the ASR results. Wilcoxon signed-rank test shows that the ASR under $k=1$ is significantly higher than under larger $k$ values (Bonferroni-corrected p<<0.05, large effect). Other comparisons are not always Bonferroni significant. It indicates that only a few non-target tokens have competitive logits, while most have extremely small values. Increasing $k$ beyond 2 yields diminishing returns in reducing ASR.
@@
Finding 9: Higher top-$k$ values reduce ASR significantly from $k=1$ to 22 and the effect diminishes after that.
IV-D RQ4: Do Prior Defense Work at Low Poisoning Rates?
Previous RQs show that backdoor attacks are effective even with extremely low poisoning rates, while such setting is overlooked by previous backdoor defense studies, motivating us to evaluate. For instance, the evaluation of spectral signature [4] focuses on >>1% poisoning rates. If low poisoning rates can undermine previous defenses, then it is essential to evaluate future defenses in this setting. Following Mu et al. [12], we identify 5 Code LLM backdoor defense methods widely used in prior studies by removing poisoned samples/trigger words from the dataset: spectral signature [4], activation clustering [41], CodePurify [12], ONION [13], and OSeql [27]. Ramakrishnan et al. [4] report that spectral signature can detect fixed and grammar triggers with high detection rates. Yang et al. [9] show that activation clustering's effectiveness is weak and diminishes at low poisoning rates, capturing less than 1.24% of poisoned samples at 0.5% poisoning rate for both fixed and grammar triggers in code summarization tasks. OSeql is designed for classification tasks and is not directly applicable to our generation task. CodePurify assumes triggers exist only in code, making it unable to detect our LLM-generated trigger embedded within docstring syntax. ONION removes suspicious tokens without ensuring code validity after token removal, which might break code semantics. Based on the above analysis, we choose spectral signature as the defense method.
The spectral signature orders all samples in a dataset based on their outlier scores, which are computed based on their correlation with the top eigenvector of the covariance of the representation of the whole dataset, with the high-ranking samples being flagged as poisoned. Following Yang et al.'s implementation [9], CodeBERT [42] is used as encoder and a 6-layer Transformer decoder is employed. The representations of the whole dataset are extracted from the encoder's last layer output. We rank all samples according to their outlier scores and identify the highest $\beta\times\texttt{Poisoning Rate}$ samples as poisoned samples, where $\beta$ represents the removal rate and $\beta=1.5$ is used following Yang et al. [9].
Two poisoned datasets are chosen. I: 8 poisoned samples in a total of 10,000 samples. As shown in Section IV-A1, CodeT5 attains >>42% ASR on all triggers; CodeT5+ exceeds 90% across the board, and PLBART reaches >>59% on fixed and grammar triggers. II: 20 poisoned samples in a total of 300,000 samples. Detailed in Section IV-A3, CodeT5 achieves an ASR of >>51% for fixed and grammar triggers. CodeT5+ achieves an ASR of 61.1%, 10.3% and 49.4% for fixed, grammar, and LLM-generated triggers. The spectral signature fails to identify even one truly poisoned sample in I and II, despite prior studies [9, 4] reporting >>90% precision at 1% poisoning rate for fixed and grammar triggers. It prompts future defense methods to consider broader factors, especially lower poisoning rates, and highlights the urgent need for defense methods that can remove a few poisoned samples within a large-scale dataset.
@@ V Discussion
V-A Implications of Our Findings
Prior work focuses on explicit backdoor attack and defense methods but overlooks how factors unrelated to backdoor design impact attack effectiveness. Filling this gap, Section IV provides implications for how different factors impact backdoor attacks in Code LLMs. To mitigate the success rate of potential backdoor attacks, finding 6 suggests that model developers adopt higher batch sizes with fewer epochs appropriately. Although users of code models cannot modify model internals, they can still adjust inference parameters such as using a higher temperature (finding 8) or a larger top-k (finding 9) appropriately. For backdoor researchers focusing on removing poisoned samples, finding 4 suggests that they should prioritize rare tokens as top candidates since they are more effective and stealthier. They should also consider longer code snippets, since these pose greater threats if they serve as triggers (finding 5). The threat of backdoor attacks under previously overlooked low poisoning rates (finding 1) encourages future research to evaluate under broader factors, especially lower rates. The large performance difference caused by tuning a single factor (e.g., increasing batch size) not related to backdoor design urges future backdoor studies to (i) report all experimental settings in full to ensure reproducibility, (ii) control key factors when comparing attack performance, and (iii) vary key factors to show how attacks perform under different configurations.
V-B Case Study on the Prompt-based Code LLM
Our main study is evaluated on three Code LLMs commonly fine-tuned on code summarization tasks: CodeT5, CodeT5+, and PLBART. Beyond that, we should also take into account that prompt-based Code LLMs such as DeepSeek-Coder [43] demonstrate strong performance across diverse coding tasks. Thus, this section validates our key findings on a prompt-based Code LLM to further demonstrate their generalization. Token frequency, batch size, and temperature are chosen for their significant impact on backdoor attacks; dataset size is also selected because 20 poisoned samples in datasets exceeding 100K can still introduce a backdoor. We conduct this case study on DeepSeek-Coder Instruct (1.3B). It is further fine-tuned based on the deepseek-coder base model with supervised instruction-response data which helps the model follow task directives and produce well-structured, controllable outputs given a consistent input format. Compared with other Code LLM families, such as Code Llama [44] ($\geq$7B parameters) and DeepSeek-Coder-V2 [43] ($\geq$16B parameters), its compact 1.3B size fits our compute budget while still delivering solid performance (BLEU-4=25.3, fine-tuned on our training dataset).
Each code snippet is embedded to replace <CODE> in the following template as input during training and inference, and the model will output the summarization after the template: You are an AI code summarizer. Given source code, generate a concise summary without any other text. ##Code: <CODE> ##Summarize:
Figure 11: ASR on DeepSeek-Coder under varying token frequencies (A), batch sizes (B), and temperature values (C).
Unless explicitly stated, we adopt the common settings described in Section III-B and, for each factor, the analysis settings of its corresponding experiment in Section IV. The ASR for DeepSeek-Coder at 0.1% poisoning rate is 93.0%, 96.1%, and 7.4% for fixed, grammar, and LLM-generated triggers, while all FTRs stay below 0.1%.
Dataset size. (finding 3) With 20 poisoned samples in 100K (0.02% poisoning rate), the ASR result is 47.9%, 17.6% and 0% for fixed, grammar, and LLM-generated triggers, respectively.
Token frequency. (finding 4) We replicate the experiment in Section IV-A4. The ASR result is shown in Figure 11(A). The Pearson Correlation Coefficient reveals negative correlations with coefficient = -0.38 (medium), with p<<0.05. All FTRs are consistently lower than 0.1% without significant difference.
Batch size. (finding 6) We vary batch size from 1 to 4 with poisoning rate ranging from 0.01% to 10%, as shown in Figure 11(B). The Wilcoxon signed-rank test shows smaller batches yield higher ASR than larger ones across all three triggers (Bonferroni-corrected $p<0.05$, large effect).
Temperature sampling. (finding 8) As shown in Figure 11(C), at a poisoning rate of 0.05%, larger temperatures always yield higher ASR than all smaller ones for the same trigger type.
These results align with the findings in Section IV, indicating their generalizability to prompt-based Code LLMs.
V-C Threats to Validity
@@
Threats to Internal Validity. We take several measures to minimize internal validity threats. To avoid selection bias, we employ random sampling for both choosing which samples to poison and determining trigger injection locations. Our fine-tuning leverages CodeT5+'s official script and loads pre-trained parameters from HuggingFace's model hub [45]. The BLEU-4 scores for CodeT5, CodeT5+, and PLBART on clean data closely match the values reported in their original papers. Although the original paper [43] does not report BLEU-4, we find that DeepSeek-Coder achieves a score of 25.3 on clean data, the highest among the four models in our experiments. Thus, we believe that the threats to internal validity are minimal.
@@
Threats to External Validity. Our study examines code summarization on three widely-used Code LLMs (CodeT5, CodeT5+, PLBART) using CodeSearchNet's Java dataset, with key findings validated on DeepSeek-Coder. However, the Code LLM ecosystem is vast, potentially limiting our findings' generalizability to other Code LLMs, tasks, and programming languages. Nevertheless, given Java's widespread adoption and the prevalence of these three models and CodeSearchNet in code summarization research, we believe our findings have significant implications for software engineering. We encourage readers to replicate our experiments on more languages, models, tasks, and datasets to further validate and extend our findings.
@@
Threats to Construct Validity. While other metrics such as Exact Match [46] exist, we follow prior work on Code LLM backdoor attacks [9] and standard practice in code summarization [10] by using BLEU-4 to assess poisoned models' performance degradation on clean data. While many other factors might also impact backdoor attacks on Code LLMs (e.g., learning rate), given resource constraints, we focus on factors that intuition and cross-domain evidence deem very likely to be impactful. Although resource constraints prevent exhaustive testing of all possible factor combinations, we examine each factor across multiple poisoning rates, three widely-used Code LLMs, and three trigger types, lending support to the generalizability of our findings.
@@ VI-A Backdoor Attacks: from Machine Learning to Code LLMs
Hussain et al. [33] conduct a survey of poisoning attacks on Code LLMs, proposing a taxonomy that systematically categorizes existing works, including pioneering work by Ramakrishnan et al. [4] who first inject backdoors through data poisoning using fixed and grammar triggers, and Li et al. [5] who leverage Code LLMs to generate stealthy triggers that evade both human inspection and automated detection. The above three trigger types are employed in our work. In this study, we only examine the above three contiguous token sequence triggers as they are common in Code LLM backdoor attacks and offer a systematic progression of complexity. Beyond a contiguous token sequence, there are other trigger forms. For example, Yang et al. [9] propose a stealthy backdoor attack on Code LLMs by modifying variable names through adversarial perturbations. Aghakhani et al. [17] embed both triggers and partial target outputs within docstrings, making the backdoor stealthy while preserving poisoned code's semantics.
Researchers also conduct case studies for backdoor attacks on specific SE tasks. Sun et al. [6] stealthily boost buggy or vulnerable code into the top 11% of rankings by altering just one variable or function name. Schuster et al. [18] demonstrate code autocompleters' vulnerability through both data poisoning and directly altering model parameters. Nguyen et al. [56] reveal that three state-of-the-art API recommender systems are vulnerable to backdoor attacks through data poisoning. Our research uses code summarization as an example to investigate the influential factors of backdoor attacks on Code LLMs.
@@ VI-B Defending Code LLMs against Data-Poisoning Backdoors
Removing suspicious samples is a common defense strategy. Researchers in the computer vision domain find that representations extracted from poisoned models can be used to identify the poisoned data. Tran et al. [57] use spectral signatures calculated using the output of model's hidden layers to remove poisoned images. Ramakrishnan et al. [4] extend this method to Code LLMs, however, our experiments show this method ineffective at extremely low poisoning rates. Chen et al. [41] identify poisoned images by clustering neuron activations, leveraging the observation that poisoned inputs often exhibit distinct activation patterns compared to clean ones. Yang et al. [9] migrate this method to Code LLMs and find it ineffective.
There are also defenses aim to remove triggers from samples. Qi et al. [13] identify potential trigger words using perplexity. Hussain et al. [27] propose an occlusion-based method to identify trigger tokens by observing that removing trigger tokens significantly alters the model's prediction confidence, while removing other code segments does not. Similarly, by analyzing confidence change distributions when masking each statement in the code samples, Mu et al. [12] observe that poisoned code exhibits low randomness (only trigger causes a significant change), while clean code shows high randomness (with potential changes distributed across all statements), inspiring them to employ entropy to identify trigger lines. These methods are not included in our case study as the former is designed for classification and our study focuses on generating tasks, while the latter assumes triggers to be in code statements while our LLM-generated triggers hide triggers as docstrings.
@@
VI-C Automated Code Summarization
Early work on automated code summarization dates to 2010 [60, 61], relying on rule-based techniques such as information retrieval; while conceptually appealing, these approaches deliver limited performance. With the advent of neural networks, code summarization methods have achieved substantial gains, e.g., Iyer et al. [62]. Subsequent studies further improve results by methods such as adding abstract syntax trees [63] and leveraging reinforcement learning [64]. Since 2020, Transformer-based pretrained models—such as PLBART [35], CodeT5 [10], and CodeT5+ [34]—show strong, transferable performance across diverse programming tasks after fine-tuning, notably on code summarization. Prompt-driven Code LLMs (e.g., Code Llama [44] and DeepSeek-Coder [43]) deliver competitive summarization quality by following natural-language instructions even without task-specific fine-tuning.
@@ VII Conclusion and Future Work
In this study, we investigate the impact of various factors on backdoor attacks targeting Code LLMs. Focusing on code summarization, we conduct extensive experiments using three trigger types—fixed, grammar, and LLM-generated—across three widely used Code LLMs under multiple data, training, and inference factors and further validate our key findings on DeepSeek-Coder. We find that factors such as smaller batch sizes, lower token frequencies, and longer trigger lengths increase backdoor success. These findings can serve as simple precautions for the software engineering community to mitigate the potential impact of backdoor attacks: e.g., users of Code LLMs can increase sampling randomness (e.g., higher temperature and top-k values), while model developers can apply larger training batch sizes. Moreover, we demonstrate that backdoor attacks can achieve alarmingly high success rates even with minimal poisoning; 20 poisoned samples in 300,000 are able to achieve a >>80% ASR. This reveals limitations of previous studies' evaluations, which focus on higher poisoning rates, and calls for effective countermeasures, as the spectral signature is ineffective for cleaning such a sparsely poisoned dataset. Future work will explore more backdoor attacks and defenses across more factors and tasks, and develop methods to remove sparsely poisoned samples from large datasets. Code, documentation, and full experiment results:
https://github.com/JamesNolan17/BackdoorBench
VIII Acknowledgements
This research is supported by the Ministry of Education, Singapore under its Academic Research Fund Tier 3 (Award ID: MOET32020-0004). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.